demand is P=400-2Q. marginal cost is 80. find the consumer surplus
1
Expert's answer
2020-05-18T11:57:34-0400

If demand is P = 400 - 2Q, then marginal revenue is:

MR = TR'(Q) = 400 - 4Q.

The profit-maximizing output is produced at MR = MC, so:

400 - 4Q = 80,

Q = 80 units.

Profit-maximizing price is:

P = 400 - 2×80 = 240.

Consumer surplus is:

CS = 0.5×(400 - 240)×80 = 6,400.